Your parents set up a trust fund for you 10 years ago that is now worth R19 671,51. If the fund earned 7% per year, how much did your parents invest?


1
Expert's answer
2022-03-23T15:56:55-0400

Present value"=\frac{FV}{(1+r)^n}"


"=\frac{R19671,51}{(1+0.07)^{10}} =R9999.99"